# Adaptive Exploration for Data-Efficient General Value Function Evaluations

### v1

@@ Abstract
General Value Functions (GVFs) (Sutton et al, 2011) are an established way to represent predictive knowledge in reinforcement learning. Each GVF computes the expected return for a given policy, based on a unique pseudo-reward. Multiple GVFs can be estimated in parallel using off-policy learning from a single stream of data, often sourced from a fixed behavior policy or pre-collected dataset. This leaves an open question: how can behavior policy be chosen for data-efficient GVF learning? To address this gap, we propose GVFExplorer, which aims at learning a behavior policy that efficiently gathers data for evaluating multiple GVFs in parallel. This behavior policy selects actions in proportion to the total variance in the return across all GVFs, reducing the number of environmental interactions. To enable accurate variance estimation, we use a recently proposed temporal-difference-style variance estimator. We prove that each behavior policy update reduces the mean squared error in the summed predictions over all GVFs. We empirically demonstrate our method’s performance in both tabular representations and nonlinear function approximation.
@@ 1 Introduction
The ability to predict many different signals is a key attribute of human, animal, and likely artificial intelligence too. In reinforcement learning (RL), a model of predictive knowledge that is grounded in an agent’s sensory experience is formulated as General Value Functions (GVFs) (Sutton et al., 2011). Each GVF is defined by a unique policy and pseudo-reward known as cumulant. GVFs generalize the classic concept of a value function. They measure an agent’s expected cumulative discounted value of cumulant, where the discount factor may vary as a function of state (instead of being fixed). A significant subclass of GVFs focus on predicting these values for a given fixed policy. These predictive GVFs can answer questions like: “How much time is there before the agent hits a wall, given its current way of moving?” or “What is the probability of the agent to collect a particular object in an allotted time, under a given policy” (White et al., 2015; Schlegel et al., 2021; Sherstan, 2020).
@@
As each GVF may be defined in terms of a unique policy-cumulant pair, this requires off-policy learning to update all GVFs in parallel. Prior works have estimated GVF values by either using a pre-collected datasets (Xu et al., 2022) or by following a fixed, randomized behavior policy (Sutton et al., 2011). However, these methods can suffer from large value estimation errors if the behavior policy or dataset is ill-suited for learning a particular GVF. Here, “ill-suited” means that the behavior policy can be very different from the GVF policies or the behavior policy does not sufficiently visit regions crucial for learning GVFs. Our work addresses this gap by focusing on the question of exploration for evaluating GVFs: how can we adapt an agent’s behavior policy to strategically sample data for multiple GVF evaluations in parallel? While exploration has been extensively studied in context of optimal control in Markov Decision Processes (MDP), the question of constructing a policy that can learn multiple quantities in parallel has remained largely untouched.
To accurately evaluate multiple GVFs in parallel, we aim to design a behavior policy that minimizes the overall mean squared error (MSE) in their predictions. An intuitive approach would be to follow each GVF’s target policy for some period of time (e.g. one episode) in a round-robin manner, and parallely update all GVFs off-policy. However, this approach can be highly data inefficient as actions are sampled according to given policies, overlooking high-variability regions crucial for learning. To overcome this limitation, we introduce an approach, GVFExplorer, that strategically directs the behavior policy to sample actions in proportion to the total variance across GVF predictions. This strategy is based on the understanding that prioritizing high-variance regions leads to more informative samples which ultimately enhance the accuracy of GVF evaluations. We use an existing temporal difference (TD) based variance estimator (Jain et al., 2021) to allow bootstrapping and iterative updates.
@@
(1) We design an adaptive behavior policy that enables learning multiple GVF predictions accurately and efficiently in parallel [Algorithm 1]. (2) We derive an iterative behavior update rule that directly minimizes the overall prediction error [Theorem 5.1]. (3) We prove that in the tabular setting, each update to the behavior policy reduces the total MSE across GVFs [Theorem 5.2]. (4) We establish the existence of a variance operator that enables us to use TD-based variance estimation [Lemma 6.1]. (5) We empirically demonstrate that GVFExplorer lowers the total MSE when estimating multiple GVFs compared to baseline approaches and enables evaluating a larger number of GVFs in parallel.
@@ 2 Related Work
In RL, the exploration literature has mainly focused on exploration for improving policy performance for a single objective (Oudeyer et al., 2007; Schmidhuber, 2010; Jaderberg et al., 2016; Machado et al., 2017; Eysenbach et al., 2018; Burda et al., 2018; Guo et al., 2022). Refer to Ladosz et al. (2022) for a detailed survey on exploration techniques in RL. While related to exploration, our work instead focuses on learning a behavior policy for evaluating multiple GVFs.
Our work is most closely related to other works on learning multiple GVFs. While Xu et al. (2022) studies a similar problem of evaluating multiple GVFs using an offline dataset, our method operates online without any pre-collected samples. Thus, our work avoids the data coverage limitations of offline approaches. Linke et al. (2020) develops exploration for learning GVFs in a stateless bandit context, which does not deal with the off-policy learning or function approximation challenges present in the full Markov Decision Process (MDP) context. Prior works like Hanna et al. (2017) learned a behavior policy for a single policy evaluation problem using a REINFORCE-style (Williams, 1992) variance-based method called BPS. This idea is similar to variance reduction techniques in Monte Carlo Simulation, which use Importance Sampling and derive a minimum-variance sampling strategy (Owen, 2013; Frank et al., 2008). Metelli et al. (2023) also extends this idea to the control setting. However, these methods are focused on single-task evaluation or control. Predicting multiple values is more challenging due to the need to carefully balance action selection among various interrelated learning problems. Perhaps the closest to our work is the algorithm of McLeod et al. (2021), which uses weight changes in the Successor Representation (SR) (Dayan, 1993) as an intrinsic reward to learn a behavior policy that supports multiple predictive tasks. GVFExplorer is simpler, optimizing the behavior policy to directly minimize total prediction error over GVFs, resulting in an intuitive variance-proportional sampling strategy. We will compare the two approaches empirically as well.
@@ 3 Preliminaries
Consider an agent interacting with the environment to obtain estimate $\mathcal{N}$ different General Value Function (GVF) (Sutton et al., 2011). We assume an episodic, discounted Markov decision process (MDP) where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the action set, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition probability function, $\Delta_{\mathcal{S}}$ is the $\mathcal{S}$-dimensional probability simplex, and $\gamma\in[0,1)$ is the discount factor.
Each GVF is conditioned on a fixed policy $\pi_{i}:\mathcal{S}\to\Delta_{\mathcal{A}}$, $i=\{1,\dots,\mathcal{N}\}$ and has a cumulant $c_{i}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. For simplicity, we assume that all cumulants are scalar, and that the GVFs share the environment discount factor $\gamma$. This eases the exposition, but our results can be extended to general multidimensional cumulants and state dependent discount factor $\gamma$.
Each GVF is a value function $V_{\pi_{i}}(s)=\mathbb{E}_{\pi_{i},\mathcal{P}}[G_{t}^{i}|s_{t}=s]$ where $G_{t}^{i}=c_{i,t}+\gamma G_{t+1}^{i}$. Each GVF can be viewed as answering the question, “what is the expected discounted sum of $c_{i}$ received while following $\pi_{i}$?” We can also define action-value GVFs: $Q_{\pi_{i}}(s,a)=c_{i}(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[V_{\pi_{i}}(s^{\prime})]$. Similar to traditional value functions, we have: $V_{\pi_{i}}(s)=\mathbb{E}_{a\sim\pi_{i}(\cdot|s)}[Q_{\pi_{i}}(s,a)]$.
At every time step $t$, the agent is in state $s_{t}$, takes an action $a_{t}$ and receives cumulant values $c_{i,t}$ for all $i\in\{1,\dots,\mathcal{N}\}$, transitioning to a new state $s_{t+1}$. This repeats until reaching a terminal state or a maximum step count. Then the agent resets to a new initial state and starts again. The agent interacts with environment using some behavior policy, $\mu:\mathcal{S}\to\Delta_{\mathcal{A}}$. The goal of the agent is to compute the approximate values $\hat{V}_{i}$ corresponding to the true GVFs value $V_{\pi_{i}}$. We formalize this objective as minimizing the Mean Squared Error (MSE) under some state weighting $d(s)$ for all GVFs:
@@
In our experiments, we use the uniform distribution for $d(s)$. In principle, this objective could be generalized to prioritize certain GVFs by using a weighted MSE.
Importance Sampling.
To estimate multiple GVFs with distinct target policies $\pi_{i}$ in parallel, off-policy learning is essential. Importance sampling (IS) is one of the primary tools for off-policy value learning (Hesterberg, 1988; Precup, 2000; Rubinstein & Kroese, 2016). IS allows to estimate the expected return $G$ under a target policy $\pi$, using trajectories from a different behavior policy $\mu$. The probability of trajectory $\tau=(s_{0},a_{0},r_{0}\dots s_{T},r_{T})$ occurring under $\pi$ is denoted by $w_{\pi}(\tau)$, and under $\mu$ by $w_{\mu}(\tau)$. The true value for any $s$ state $V_{\pi}(s)=\mathbb{E}_{\tau\sim\pi}[G_{t}|s_{t},a\sim\pi]$ can be estimated by sampling actions under $\mu$ as $\hat{V}_{\pi}^{\mu}(s)=\dfrac{1}{K}\sum_{j=1}^{K}\dfrac{w_{\pi}(\tau_{j})}{w_{\mu}(\tau_{j})}\{G_{t}|s_{t}=s,\tau_{j}\sim\mu\}$. This IS estimate is unbiased, i.e. $\mathbb{E}_{\tau\sim\mu}[\hat{V}_{\pi}^{\mu}(s)]=V_{\pi}(s)$. There are also variations of IS that provides lower variance, such as per-decision IS, weighted IS and discounted stationary distribution IS (Precup, 2000).
4 Problem Formulation
As described in the previous section, the goal of the agent is to minimize the total mean squared error (MSE) across the given GVFs (Equation 1). Note that $\text{MSE}=\text{Variance}+\text{Bias}^{2}$. As we will be using unbiased IS estimation for off-policy correction, the task of reducing the total MSE essentially becomes minimizing the total variance across GVFs. Thus, the crux of our problem is to find a behavior policy that gathers data in such a way as to minimize the variance across the set of GVFs.
Prior works (Kahn & Marshall, 1953; Owen, 2013) have studied a simpler version of the above problem, seeking an optimal behavior policy $\mu^{*}$ for estimating a single target policy value using IS through variance reduction:
$\min_{\mu}\text{Var}_{a\sim\mu}\left(\frac{w_{\pi}(\tau)}{w_{\mu}(\tau)}G(\tau)\right)\implies w_{\mu^{*}}(\tau)=\dfrac{w_{\pi}(\tau)}{\mathbb{E}_{\tau\sim\pi}[G(\tau)]}G(\tau).$
However, this analytical solution of $\mu^{*}$ is infeasible for estimating multiple GVFs due to its dependence on the unknown quantity $\mathbb{E}_{\tau\sim\pi}[G(\tau)]=V_{\pi}$. To address this limitation and estimate multiple GVFs as well, our approach involves optimizing $\mu^{*}$ to minimize the total variance in cumulants across all GVFs, under state weighting $d(s)$:
$\mu^{*}=\arg\min_{\mu}\sum_{i=1}^{\mathcal{N}}\sum_{s\in\mathcal{S}}d(s)\text{Var}_{\pi_{i}}(G_{t}|s_{t}=s,a\sim\mu).$ (2)
@@ 4.1 Objective Function
To solve this optimization problem, we will rely on the variance function $\boldsymbol{M_{\pi}^{\mu}(s)}$, which measures the variance in the return of target policy $\pi$ in a state $s$ when actions are sampled under a different behavior policy $\mu$. In-depth insights into the variance operator is explored in Section 6. The variance function for a given state is defined as:
$M_{\pi}^{\mu}(s)=\text{Var}_{\pi}(G_{t}|s_{t}=s,a\sim\mu)$
Similarly, for each state-action pair, we can define:
$M_{\pi}^{\mu}(s,a)=\text{Var}_{\pi}(G_{t}|s_{t}=s,a_{t}=a,a^{\prime}\sim\mu)$
In this work, the objective is to find an optimal behavior policy $\mu^{*}$ that efficiently collects data to minimize the sum of variances $M_{\pi_{\{1\dots\mathcal{N}\}}}$ under some desired state distribution $d(s)$. Thus, the objective is:
@@
Next, we present an iterative update for the behavior policy obtained by solving the objective in Equation 3.
5 Theoretical Framework

### v2

@@ Abstract
General Value Functions (GVFs) (Sutton et al., 2011) represent predictive knowledge in reinforcement learning. Each GVF computes the expected return for a given policy, based on a unique reward. Existing methods relying on fixed behavior policies or pre-collected data often face data efficiency issues when learning multiple GVFs in parallel using off-policy methods. To address this, we introduce GVFExplorer, which adaptively learns a single behavior policy that efficiently collects data for evaluating multiple GVFs in parallel. Our method optimizes the behavior policy by minimizing the total variance in return across GVFs, thereby reducing the required environmental interactions. We use an existing temporal-difference-style variance estimator to approximate the return variance. We prove that each behavior policy update decreases the overall mean squared error in GVF predictions. We empirically show our method’s performance in tabular and nonlinear function approximation settings, including Mujoco environments, with stationary and non-stationary reward signals, optimizing data usage and reducing prediction errors across multiple GVFs.
@@ 1 Introduction
The ability to make multiple predictions is a key attribute of human, animal, and artificial intelligence. Sutton et al. (2011) introduced General Value Functions (GVFs) which consists of several independent sub-agents, each responsible for answering specific predictive knowledge about the environment. Each GVF consists of a unique - policy, custom reward function called cumulant and state-dependent discount factor - to calculate the cumulative discounted cumulant. For example, a GVF can predict the expected number of times an agent will bump into the wall under a given policy (White et al., 2015; Schlegel et al., 2021; Sherstan, 2020). In essence, GVFs generalizes the standard value function to address a wider range of predictive questions, making them a powerful tool for intelligent systems.
@@
Prior works have used either a fixed randomized behavior policy (Sutton et al., 2011) or pre-collected datasets (Xu et al., 2022) to update all GVFs in parallel using off-policy learning. However, these methods can result in large value estimation errors if the behavior policy significantly diverges from the GVF policies. Our work addresses this gap by focusing on the question of exploration for evaluating GVFs: how can we adapt an agent’s behavior policy to data-efficiently sample for evaluating multiple GVFs in parallel? While exploration has been extensively studied in context of optimal control in Markov Decision Processes (MDP), the question of constructing a policy that can learn multiple quantities in parallel has remained largely untouched.
To accurately evaluate multiple GVFs in parallel, we aim to design a behavior policy that minimizes the overall mean squared error (MSE) in their predictions. A natural approach might involve following each GVF’s target policy for some period of time (e.g. one episode) in a round-robin manner while concurrently updating all GVFs off-policy. However, this approach can be highly data inefficient as actions are sampled according to given policies, potentially overlooking actions from states where expected return is uncertain. To achieve better value estimation in fewer samples, it is essential to focus on state-action pairs with high variance in return, as these pairs would exhibit greater uncertainty in their mean return. Therefore, a behavior policy should visit such pairs more frequently to offset higher variance in return. Consider an analogy of two-arm bandit problem: fewer samples are needed for a constant reward arm for accurate value estimation, whereas an arm with a variable reward demands a greater number of samples to achieve the same level of certainty. We empirically support this claim by comparing round-robin and our proposed approach later.
With this motivation, we introduce GVFExplorer, that adaptively learns a behavior policy which minimize the total MSE across all GVF predictions. GVFExplorer leverages the existing off-policy temporal difference (TD) based estimator of variance in return distribution (Sherstan et al., 2018; Jain et al., 2021) to guide the behavior policy. The strategy is to frequently take actions that might have more unpredictable outcomes (high variance in return). By sampling them more, agent can estimate the mean return better with fewer interactions, thus effectively lowering the overall MSE in the GVF predictions.
GVFExplorer optimizes the data usage and reduces the prediction error, offering a scalable solution for complex environments. This is particularly valuable for real-world applications like personalized recommender systems (Parapar & Radlinski, 2021; Tang et al., 2015), where it can enable efficient evaluation of personalized policies based on diverse user preferences (reward functions) (Li et al., 2024), leveraging shared knowledge for improved accuracy.
@@
(1)We design an adaptive behavior policy that enables accurate and efficient learning of multiple GVF predictions in parallel [Algorithm 1]. (2) We derive an iterative behavior update rule that directly minimizes the overall prediction error [Theorem 4.1]. (3) We prove in tabular setting that each iterative update to the behavior policy causes the total MSE across GVFs to be less than equal to one from the old policy [Theorem 4.2]. (4) We establish the existence of a variance operator that enables us to use TD-based variance estimation [Lemma 5.1]. (5) We empirically demonstrate in both tabular and Mujoco environments that GVFExplorer lowers the total MSE when estimating multiple GVFs compared to baseline approaches and enables evaluating a larger number of GVFs in parallel.
@@ 2 Related Work
Exploration in reinforcement learning (RL) has predominantly focused on improving policy performance for a single objective (Oudeyer et al., 2007; Schmidhuber, 2010; Jaderberg et al., 2016; Machado et al., 2017; Eysenbach et al., 2018; Burda et al., 2018; Guo et al., 2022). Refer to Ladosz et al. (2022) for a detailed survey on exploration techniques in RL. While related to exploration, these works differ from ours, as they concentrate on optimizing policies for single objective rather than evaluating multiple GVFs (policy-cumulant pair) simultaneously.
Our work is most closely related to other works on learning multiple GVFs. Xu et al. (2022) address a similar problem by evaluating multiple GVFs using an offline dataset, but our method operates online, avoiding the data coverage limitations of offline approaches. Linke et al. (2020) develops exploration strategies for GVFs in a stateless bandit context, which does not deal with the off-policy learning or function approximation challenges present in the full Markov Decision Process (MDP) context. In a single bandit problem, Antos et al. (2008); Carpentier et al. (2015), show that the optimal data collection strategy to estimate mean rewards of arms is to sample proportional to each arm’s variance in reward. Prior works like Hanna et al. (2017) learned a behavior policy for a single policy evaluation problem using a REINFORCE-style (Williams, 1992) variance-based method called BPS. This idea extends on the similar principles of using Importance Sampling in Monte Carlo simulations for finding optimal sampling policy based on variance minimization  (Owen, 2013; Frank et al., 2008). Metelli et al. (2023) extends this idea to the control setting. However, these methods are limited to single-task evaluation or control. Evaluating multiple policies simultaneously is more complex, requiring careful balance in action selection among interrelated learning problems. Perhaps the closest work to ours is by McLeod et al. (2021), which uses the changes in the weights of Successor Representation (SR) (Dayan, 1993) as an intrinsic reward to learn a behavior policy that supports multiple predictive tasks. GVFExplorer approach is simpler, as it directly optimizes the behavior policy to minimize the total prediction error over GVFs, resulting in an intuitive variance-proportional sampling algorithm. We will compare the two approaches empirically as well.
@@ 3 Preliminaries
Consider an agent interacting with the environment to obtain estimates of $\mathcal{N}$ different General Value Function (GVF) (Sutton et al., 2011). We assume an episodic, discounted Markov decision process (MDP) where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the action set, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition probability function, $\Delta_{\mathcal{S}}$ is the $|\mathcal{S}|$-dimensional probability simplex, and $\gamma\in[0,1)$ is the discount factor.
Each GVF is conditioned on a fixed policy $\pi_{i}:\mathcal{S}\to\Delta_{\mathcal{A}}$, $i=\{1,\dots,\mathcal{N}\}$ and has a cumulant $c_{i}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. For simplicity, we assume that all cumulants are scalar, and that the GVFs share the environment discount factor $\gamma$. This eases the exposition, but our results can be extended to general multidimensional cumulants and state dependent discount factor. Each GVF is a value function $V_{\pi_{i}}(s)=\mathbb{E}_{\pi_{i},\mathcal{P}}[G_{t}^{i}|s_{t}=s]$ where $G_{t}^{i}=c_{i,t}+\gamma G_{t+1}^{i}$. Each GVF can be viewed as answering the question, “what is the expected discounted sum of $c_{i}$ received while following $\pi_{i}$?” We can also define action-value GVFs: $Q_{\pi_{i}}(s,a)=c_{i}(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[V_{\pi_{i}}(s^{\prime})]$, with $V_{\pi_{i}}(s)=\mathbb{E}_{a\sim\pi_{i}(\cdot|s)}[Q_{\pi_{i}}(s,a)]$.
At each time step $t$, the agent in state $s_{t}$, takes an action $a_{t}$ and receives cumulant values $c_{i,t}$ for all $i\in\{1,\dots,\mathcal{N}\}$, transitioning to a new state $s_{t+1}$. This repeats until reaching a terminal state or a maximum step count. Then the agent resets to a new initial state and starts again. The agent interacts with environment using a behavior policy, $\mu:\mathcal{S}\to\Delta_{\mathcal{A}}$. The goal is to approximate values $\hat{V}_{i}$ corresponding to the true GVFs value $V_{\pi_{i}}$. We formalize the objective as minimizing the Mean Squared Error (MSE) under some state weighting $d(s)$ for all GVFs:
@@
In our experiments, we use the uniform distribution for $d(s)$. This objective can be generalized to prioritize certain GVFs using a weighted MSE.
Importance Sampling (IS).
To estimate multiple GVFs with distinct target policies $\pi_{i}$ in parallel, off-policy learning is essential. Importance sampling (IS) is one of the primary tools for off-policy value learning (Hesterberg, 1988; Precup, 2000; Rubinstein & Kroese, 2016), allowing estimation of value function under target policy $\pi$ using samples from different behavior policy $\mu$. In context of off-policy Temporal Difference (TD) learning (Sutton & Barto, 2018), IS ratio, $\rho_{t}=\frac{\pi(a_{t}|s_{t})}{\mu(a_{t}|s_{t})}$, is used to adjust the updates to ensure unbiased value estimates. The update rule is given as $\hat{Q}(s_{t},a_{t})=\hat{Q}(s_{t},a_{t})+\alpha\left(c_{t}+\gamma\rho_{t+1}\hat{Q}(s_{t+1},a_{t+1})-\hat{Q}(s_{t},a_{t})\right),$ where $\alpha$ is the learning rate. This update rule ensures that estimated value function $\hat{Q}$ converges to correct value $Q_{\pi}$ under policy $\pi$, despite the samples being generated from a behavior policy $\mu$.
4 Behavior Policy Optimization
As described in the previous section, the goal of the agent is to minimize the total mean squared error (MSE) across the given GVFs (Eq. 1). Note that $\text{MSE}=\text{Variance}+\text{Bias}^{2}$. For the algorithm’s derivation, we will use unbiased IS estimation for off-policy correction, which shifts the task of minimizing MSE to reducing the total variance across GVFs. Thus, the core problem is to design a behavior policy that collects data to minimize the variance in return across GVFs and accurately estimate multiple GVF value functions.
The problem of estimating a single target policy’s value $V_{\pi}$ is well studied in the literature. Kahn & Marshall (1953); Owen (2013) uses IS through variance minimization to find the optimal behavior policy $\mu^{*}$. Owen (2013) also showed that $\mu^{*}$ can lead to improved performance by using off-policy value estimation over direct sampling from $\pi$, with the improvement directly related to the degree of variance reduction. However, their analytical solution of obtaining $\mu^{*}$ depends on an unknown quantity $V_{\pi}$, making it impractical. Nonetheless, the insight suggests that minimizing variance may also enhance performance in policy control for multiple GVF scenarios. In this work, however, we focus solely on policy evaluation, laying the foundation for potential future extensions to policy control.
@@ 4.1 Objective Function
We propose GVFExplorer to address the above limitation and extend the problem to accurately estimate multiple GVF values. GVFExplorer takes as input the GVF target policies $\pi_{i=\{1,\dots,\mathcal{N}\}}$, uses off-policy algorithm to sample the data from behavior policy $\mu$ (which is optimized over time) and outputs the GVF estimates $\hat{V}_{i=\{1,\dots,\mathcal{N}\}}$. Since our objective is to find a behavior policy that minimizes the variance in return across multiple GVFs, we use existing off-policy TD-style variance estimator (Sherstan et al., 2018). This estimator allows to bootstrap the target values and iteratively update the variance function, making the solution scalable to complex domains.
We define the variance function by $\boldsymbol{M_{\pi}^{\mu}(s)}$, which measures the variance in the return of target policy $\pi$ in a given state $s$ when actions are sampled under a different behavior policy $\mu$. In-depth insights into the variance operator is explored in Sec. 5. The variance function for a given state and a give state-action pair is defined respectively:
$M_{\pi}^{\mu}(s)=\text{Var}_{\pi}(G_{t}|s_{t}=s,a\sim\mu)\quad\text{and}\quad M_{\pi}^{\mu}(s,a)=\text{Var}_{\pi}(G_{t}|s_{t}=s,a_{t}=a,a^{\prime}\sim\mu).$
Our objective is to find an optimal behavior policy $\mu^{*}$ that efficiently collects a single stream of experience to minimize the sum of variances $M_{\pi_{\{1\dots\mathcal{N}\}}}$ under some state distribution $d(s)$, as,
@@
We solve the above objective function iteratively. At each iteration $k$, GVFExplorer produces a behavior policy $\mu_{k}$. The behavior policy interacts with the environment and gathers data. Using this data, any off-policy TD algorithm can be used to iteratively estimate the variance function $M_{\pi}^{\mu_{k}}$. This variance function is used to update to a better policy $\mu_{k+1}$, continuing the cycle similar to policy iteration in standard RL.
Our aim is to iteratively improve behavior policy and decrease variance functions to estimate the GVF values $V_{\pi_{i=\{1,2,\dots,\mathcal{N}\}}}$ with reducing MSE:
[…]
4.2 Theoretical Solution